OpenStudy (anonymous):

Solve the volume of a cone formula, V=1/3 πr^2 h, for h

OpenStudy (alexandervonhumboldt2):

v*3=pir^2h v*3/pi/r^2=h

OpenStudy (alexandervonhumboldt2):

h=v*3/pi/r^2